# Constructing Safety Cases for AI Systems: A Reusable Template Framework

### v1

@@ 4.5 Evidence taxonomy
The taxonomy in Figure 7 organises AI safety evidence into seven primary categories including empirical, comparative, risk-based, expert-derived, formal methids, operational, and mechanistic. Each represents distinct modes of substantiation that map to different reasoning logics and argument functions introduced in Section 4.4.
@@
The quality focus of empirical evidence is on representativeness of datasets, reproducibility of test procedures, and coverage of operational contexts.
@@
Comparative benchmarking evidence demonstrates relative safety through reference evaluations. It strongly supports comparative arguments. The evidence type includes benchmarking results, comparative performance on safety-related benchmarks or competitions are used as controlled studies comparing the AI system to human or legacy systems using equivalent metrics.
The quality focus includes fairness of comparators, equivalence of evaluation scope, and statistical significance of differences.
@@
  • Expert assessment is formal elicitation or review by qualified experts estimating likelihood, severity, or adequacy of controls. It includes Delphi studies.
@@
  • Forecasting and horizon-scanning evidence is anticipatory analyses identifying emergent capabilities, threat vectors, or long-term socio-technical impacts.
The key quality aspects include diversity of expertise, transparency of reasoning, and explicit representation of uncertainty and consensus.
@@
Formal verification provides mathematically rigorous assurance that specified safety or containment properties hold. Typical methods include model checking, theorem proving, static analysis, and symbolic execution to prove invariants, enforceability of rules, and bounded-autonomy constraints. It primarily supports deductive reasoning within demonstrative and capability-oriented arguments.
The quality focus is on the soundness of proofs, tool qualification, and correspondence between the verified model and the implemented system, with clear assumption traceability and coverage of critical properties.
@@
Operational evidence validates safety during real-world use. It supports demonstrative and statistical reasoning arguments by confirming that implemented controls remain effective.
@@
The quality focus is on recency of data, authenticity of records, and traceability to system versions.
@@
The evidence type includes interpretability analysis such as attribution mapping, feature tracing, or causal-path exploration explaining why behaviour is safe or unsafe. It also includes formal verification such as proofs or static analyses demonstrating that safety or containment properties hold within subsystems.
Evidence must also satisfy quality criteria such as independence, coverage, recency, reproducibility, and representativeness, so that its credibility can be systematically assessed.
4.6 Composability and reusability of the framework
The true power of these templates lies in their composability. A full safety case can be built as a network of reusable modules. Each module is reusable across systems and contexts, with only the specific claim instantiation and evidence artefacts substituted. The result is a structured but flexible assurance framework; one capable of supporting AI systems whose safety can only be demonstrated through continual evaluation and rediscovery.
@@ 5 Safety Case Patterns for AI Systems
In this section, we introduce four representative patterns correspond to the unique characteristics of modern AI assurance: "discovery-driven evaluation", "marginal risk without ground truth", "continuous system evolution", and "threshold-based acceptability decisions". We also discuss integration of these patterns into composite safety cases used in practice.
@@ 5.1 Discovery-driven evaluation pattern
Traditional engineering safety cases rely on full hazard enumeration before operation. For AI systems, this is infeasible: new capabilities and risks are discovered only after deployment-like testing. The assurance challenge is to justify safety amid incomplete knowledge and continuous empirical discovery.
@@
  • Marginal safety claims comparing to a reference system or control scenario.
@@ 5.2 Marginal-risk pattern without ground truth
This pattern operationalises marginal safety reasoning when ground truth is absent. It leverages multi-dimensional proxies such as predictability, capability, and interactive robustness, to provide a defensible, quantitative and qualitative basis for comparing an AI system’s safety against human or legacy baselines. It enables deployment decisions grounded in relative rather than absolute assurance, while preserving transparency in the reasoning chain and quality of evidence.
@@ 5.3 Continuous-evolution pattern
Problem AI systems undergo constant modification such as model updates, retraining, data refreshes, scaffolding updates, or integration with new tools. Each change potentially alters risk characteristics, invalidating static evidence. The assurance challenge is to maintain a living safety case that stays valid through system evolution.
@@ 5.4 Threshold-comparator pattern
  • Constrained-based risk claims with explicit numerical limits.
@@ 5.5 Integrating patterns into composite safety cases
Reusable patterns therefore enable composable assurance, the ability to build, maintain, and audit AI safety cases as evolving mosaics rather than static documents. This compositional approach allows alignment with both rapid development cycles and dynamic regulatory expectations.
@@ 6.1 Case system overview and context
The pilot AI-based tender evaluation system replaces one of the two human reviewers with an AI model. The AI system was deliberately not trained on historical data. This design choice was critical to avoid bias and artificial correlation in the evaluation process. Training on historical samples would have caused the “agreement” metrics to become inflated, as outcomes would be partially determined by prior examples rather than genuine reasoning. Instead, the system relied on a prompt-based approach using GPT-4o, where both the document under evaluation and the criteria document (e.g., policy documents and scoring rubrics) were supplied at inference time. The system assists the remaining human reviewer by producing structured assessments and justifications. Its purpose is to improve efficiency and consistency while maintaining or enhancing fairness, transparency, and procedural integrity.
A key assurance challenge is the absence of an absolute ground truth: there are no definitive “correct” tender scores, only prior human judgements that vary across reviewers and contexts. Hence, safety and reliability must be justified marginally, that the new AI + human configuration performs no worse (and ideally better) than the existing human + human process across comparable criteria.
@@ 6.2 Applied pattern: Marginal-Risk Pattern without Ground Truth
  • Normative (Deductive reasoning) supports assurance by referencing accepted practice, policy, or precedent, establishing that the comparator system or metric selection is valid. In this case, the comparator (human-human baseline) is justified as an appropriate reference based on regulatory and procedural acceptance within public-sector procurement.
@@
Comparative Inductive Yes Selected as primary argument type for this case. Suitable when a validated human baseline exists, enabling empirical “no-worse-than” reasoning across fairness, consistency, and interpretability metrics. Comparative evaluation results, expert review reports, red-team findings, qualitative disagreement analysis.
Risk-based (Threshold-based) Statistical Yes Selected as as complementary argument function to express quantitative confidence. Aggregates multiple proxy metrics and demonstrates non-inferiority ($\Delta\leq\delta$) with statistical confidence level ($\geq\alpha$). Aggregated performance metrics, confidence intervals, bootstrapped significance tests, uncertainty propagation analysis.
Normative Deductive No Not selected as a primary argument in this case. The comparator (human-human baseline) and associated procurement policies are already formally accepted within governance processes; therefore, no additional normative justification is required. (If applied) Regulatory standards, policy guidelines, procurement framework documentation.
@@
To support the comparative argument (e.g., decision consistency), Threshold-based argument is defined as follows.
@@
The threshold-based argument provides quantitative reinforcement to the comparative argument by statistically verifying that observed differences between the AI–human and human–human evaluations remain within the acceptable tolerance range. It translates the qualitative claim of “no-worse-than” performance into measurable confidence bounds, thereby strengthening the empirical credibility of the primary argument.
@@
As mentioned earlier, there is no definitive ground truth in this case. Accordingly, the evaluation metrics in this case follow the principles outlined in Section 5, relying on predictability-based measures such as consistency and consensus, rather than absolute correctness. These metrics operationalise the “no-ground-truth” condition by quantifying relative coherence between human and AI reviewers.
To support the comparative and threshold-based arguments, Comparative benchmarking evidence was collected from controlled evaluation trials involving 200 synthetic tender cases.
The following statement clearly explains what assurance activities were performed, summarises what tangible artefacts were produced or collected, and links those artefacts explicitly to the argument they support.
@@
    The differences (marginal risk: e.g., -0.2% gap of inconsistency rate) clearly show that the new system (AI–human) is no worse (slightly better) than exisitng system (human-human).
    The differences in decision inconsistency between the AI–human and human–human configurations ($\Delta R=-0.2\%$) shows that the new system performs no worse and slightly better than the existing process.
@@
The findings indicate that the AI-based review process can be safely adopted alongside or as a partial replacement for the existing human-human review workflow. Adoption is expected to reduce reviewer workload and resource demand, while maintaining consistency, fairness, and transparency in decision-making. Continuous monitoring of agreement and fairness indicators is recommended to ensure that performance remains within the validated acceptance threshold.
@@ 7 Limitations and Future Work
This framework aims to establish reusable, composable templates for AI safety cases that reflect the empirical and evolving nature of AI systems. Several limitations remain.
Scalability of dynamic safety cases
Dynamic safety cases require continuous evidence ingestion, verification, and version tracking. For continuously evolving systems, achieving through-life safety assurance is challenging due to the fundamental mismatch between rapid, non-deterministic AI evolution and the need for rigorous assurance documentation [S50-dynamic-assurance, S51-dynamic-safety-cps, S86-dynamic-open-systems]. Key limitations include continuous evidence overload from real-time operational data, the computational complexity of automatically updating argument structures when AI behaviour changes, and the lack of unified standards and robust automation infrastructure for managing dynamic evidence and argument revision. Implementing them at the scale of frontier AI systems will require automation infrastructure, tooling standards, and agreement on trusted data pipelines for live safety indicators.
Interoperability across regulatory regimes
Different jurisdictions emphasise distinct aspects of AI safety, capability thresholds, model access control, or human oversight. Ensuring that reusable templates can be parameterised for diverse regulatory contexts without fragmentation is challenging due to divergent regulatory (e.g., EU AI Act [act2024eu]) philosophies, varied evidential standards and levels of assurance across sectors [S112-certification-regulation], and the difficulty of translating high-level legal requirements into structured, auditable technical claims within CAE templates [S9-gsn-ai-regulation]. Achieving effective interoperability requires developing modular template components that can be selectively activated and customized based on targeted jurisdiction and required assurance level.
Evaluation reproducibility
@@
Relationship to verification processes
This paper treats verification outcomes (e.g., test results, certificates, proofs, or monitoring summaries) as evidence inputs to an AI safety case, but does not model the verification processes that produce them. In particular, we do not characterise how behavioural properties are generated, checked, or enforced at design time or runtime. These concerns are increasingly addressed by verifiability-first AI engineering approaches [zhuverifiability], which focus on structuring systems so that behavioural claims can be verified at scale. The contribution of this paper is complementary. It focuses on how verified artefacts, once available, are assembled, contextualised, and justified within a structured safety argument.
Future directions
Further research should operationalise the template framework into a tool-supported Safety Case Pattern Library, enabling instantiation and continuous management across AI lifecycles. Key directions include formalising uncertainty propagation between reasoning modes, developing composite safety metrics for dynamic systems, and standardising comparator-based marginal-risk evaluations, all representing promising steps for both scientific and regulatory progress. In addition, three priorities emerge for our next step work: First, using LLMs to automatically generate AI safety cases or CAE templates could significantly support dynamic safety case construction. Recent studies (e.g., S14,S51,S54,S74) show that LLMs can create the structural parts of safety case elements (e.g., CAE) and speed up early safety documentation, though they have difficulty with accuracy and traceable evidence without human checking. This points toward a hybrid human-AI assurance paradigm where LLMs serve as drafting accelerators while humans validate the results. Second, integrating AI risk management and assessment frameworks [NOPSEMA2025safetycase] and operational data, (e.g., guardrail logs [shamsujjoha_ICSA-2025], evaluation results [S116-evaluation-ai-scheming], and AgentOps observability data [dong2024agentops1]) across AI safety case pipeline, could provide a more comprehensive evidence repository to support the full lifecycle of AI safety case construction, verification, and governance (as illustrated in Figure 3 ecosystem pipeline). Third, developing systematic methods for mapping broad legal requirements into specific, measurable technical claims using CAE templates (e.g., S9) would bridge the gap between regulatory mandates and technical assurance. These directions represent critical steps toward making rigorous, evidence-based safety assurance practical and scalable for frontier AI systems.
@@ 8 Conclusion
In this study, we presented a reusable safety-case template framework for AI systems. The framework uses a clear Claims–Arguments–Evidence (CAE) structure. It also provides AI-specific taxonomies for claim types, argument forms, and evidence families. These elements help practitioners build safety arguments in a consistent and reviewable way. Our proposed templates address recurring assurance problems that appear in many AI deployments. We include end-to-end patterns for evaluation without ground truth, managing dynamic model updates, and making threshold-based risk decisions. This extends safety-case practice to areas where traditional safety cases often do not fit well. It also supports safety cases that can be checked and updated as the system changes.
In addition, we embedded the templates into a continuous assurance pipeline. This pipeline links safety claims to live metrics and governance artifacts for ongoing monitoring. We applied the approach in a real-world case study on a government AI-based tender evaluation system. The case study shows how the templates can be used in practice and what evidence can be collected to support key claims. Overall, this template-based framework provides a reusable foundation for AI safety cases that must evolve over time. However, it is not a complete solution. Key remaining challenges include scaling dynamic safety cases, supporting different regulatory settings, and improving evaluation reproducibility.

### v3

@@ 4.5 Evidence
The taxonomy in Figure 7 organises AI safety evidence into seven primary categories, including empirical, comparative, risk-based, expert-derived, formal methods, operational, and mechanistic. Each represents distinct modes of substantiation that map to different reasoning logics and argument functions introduced in Section 4.4.
@@
The quality focus of empirical evidence is on the representativeness of datasets, the reproducibility of test procedures, and the coverage of operational contexts.
@@
Comparative benchmarking evidence demonstrates relative safety through reference evaluations. It strongly supports comparative arguments. The evidence type includes benchmarking results, comparative performance on safety-related benchmarks, or competitions are used as controlled studies comparing the AI system to human or legacy systems using equivalent metrics.
The quality focus includes the fairness of comparators, the equivalence of the evaluation scope, and the statistical significance of differences.
@@
  • Expert assessment is formal elicitation or review by qualified experts estimating the likelihood, severity, or adequacy of controls. It includes Delphi studies.
@@
  • Forecasting and horizon-scanning evidence is anticipatory analysis that identifies emergent capabilities, threat vectors, or long-term socio-technical impacts.
The key quality aspects include the diversity of expertise, the transparency of reasoning, and the explicit representation of uncertainty and consensus.
@@
Formal verification provides mathematically rigorous assurance that specified safety or containment properties hold. Typical methods include model checking, theorem proving, static analysis, and symbolic execution to establish invariants, enforce rules, and verify bounded-autonomy constraints. It primarily supports deductive reasoning within demonstrative and capability-oriented arguments.
The quality focus is on the soundness of proofs, tool qualification, and the correspondence between the verified model and the implemented system, with clear traceability of assumptions and coverage of critical properties.
@@
Operational evidence validates safety during real-world use. It supports both demonstrative and statistical arguments by confirming that the implemented controls remain effective.
@@
The quality focus is on data recency, record authenticity, and traceability to system versions.
@@
The evidence type includes interpretability analysis, such as attribution mapping, feature tracing, or causal-path exploration, that explains why behaviour is safe or unsafe. It also includes formal verification, such as proofs or static analyses, demonstrating that safety or containment properties hold within subsystems.
Evidence must also satisfy quality criteria such as independence, coverage, recency, reproducibility, and representativeness, enabling its credibility to be assessed systematically.
4.6 Composability and reusability of the template
The true power of these templates lies in their composability. A full safety case can be built as a network of reusable modules. Each module is reusable across systems and contexts, with only the specific claim instantiation and evidence artefacts substituted. The result is a structured but flexible assurance design; one capable of supporting AI systems whose safety can only be demonstrated through continual evaluation and rediscovery.
@@ 5 Safety Case Patterns for AI Systems
In this section, we introduce four representative patterns that correspond to the unique characteristics of modern AI assurance: "discovery-driven evaluation", "marginal risk without ground truth", "continuous system evolution", and "threshold-based acceptability decisions". We also discuss integrating these patterns into composite safety cases used in practice.
@@ 5.1 Discovery-driven evaluation pattern
Traditional engineering safety cases rely on a full hazard enumeration prior to operation. For AI systems, this is infeasible: new capabilities and risks are discovered only after deployment-like testing. The assurance challenge is to justify safety in the face of incomplete knowledge and ongoing empirical discovery.
@@
  • Marginal safety claims compared to a reference system or control scenario.
@@ 5.2 Marginal-risk pattern without ground truth
This pattern operationalises marginal safety reasoning when ground truth is absent. It leverages multi-dimensional proxies such as predictability, capability, and interactive robustness to provide a defensible, quantitative, and qualitative basis for comparing an AI system’s safety against human or legacy baselines. It enables deployment decisions grounded in relative rather than absolute assurance, while preserving transparency in the reasoning chain and quality of evidence.
@@ 5.3 Continuous-evolution pattern
Problem AI systems undergo continual modification, including model updates, retraining, data refreshes, scaffolding updates, and integration with new tools. Each change potentially alters risk characteristics, invalidating static evidence. The assurance challenge is to maintain a living safety case that stays valid through system evolution.
@@ 5.4 Threshold-comparator pattern
  • Constraint-based risk claims with explicit numerical limits.
@@ 5.5 Integrating patterns into composite safety cases
Reusable patterns, therefore, enable composable assurance, the ability to build, maintain, and audit AI safety cases as evolving mosaics rather than static documents. This compositional approach allows alignment with both rapid development cycles and dynamic regulatory expectations.
@@ 6.1 Case system overview and context
The pilot AI-based tender evaluation system replaces one of the two human reviewers with an AI model. The AI system was deliberately not trained on historical data. This design choice was critical to avoid bias and artificial correlation in the evaluation process. Training on historical samples would have inflated the “agreement” metrics, as outcomes would have been partially determined by prior examples rather than genuine reasoning. Instead, the system relied on a prompt-based approach using GPT-4o, in which both the document under evaluation and the criteria document (e.g., policy documents and scoring rubrics) were provided at inference time. The system assists the remaining human reviewer by producing structured assessments and justifications. Its purpose is to improve efficiency and consistency while maintaining or enhancing fairness, transparency, and procedural integrity.
A key assurance challenge is the absence of an absolute ground truth: there are no definitive “correct” tender scores, only prior human judgements that vary across reviewers and contexts. Hence, safety and reliability must be justified marginally, so that the new AI + human configuration performs no worse (and ideally better) than the existing human + human process across comparable criteria.
@@ 6.2 Applied pattern: Marginal-Risk Pattern without Ground Truth
  • Normative (Deductive reasoning) supports assurance by referencing accepted practice, policy, or precedent, establishing that the comparator system or metric selection is valid. In this case, the comparator (human-human baseline) is justified as an appropriate reference, given regulatory and procedural acceptance in public-sector procurement.
@@
Comparative Inductive Yes Selected as the primary argument type for this case. Suitable when a validated human baseline exists, enabling empirical “no-worse-than” reasoning across fairness, consistency, and interpretability metrics. Comparative evaluation results, expert review reports, red-team findings, and qualitative disagreement analysis.
Risk-based (Threshold-based) Statistical Yes Selected as a complementary argument function to express quantitative confidence. Aggregates multiple proxy metrics and demonstrates non-inferiority ($\Delta\leq\delta$) with statistical confidence level ($\geq\alpha$). Aggregated performance metrics, confidence intervals, bootstrapped significance tests, and uncertainty propagation analysis.
Normative Deductive No Not selected as a primary argument in this case. The comparator (human-human baseline) and the associated procurement policies have already been formally accepted within governance processes; therefore, no additional normative justification is required. (If applied) Regulatory standards, policy guidelines, and procurement framework documentation.
@@
To support the comparative argument (e.g., decision consistency), the threshold-based argument is defined as follows.
@@
The threshold-based argument provides quantitative support for the comparative argument by statistically verifying that observed differences between AI–human and human–human evaluations remain within an acceptable tolerance range. It translates the qualitative claim of “no-worse-than” performance into measurable confidence bounds, thereby strengthening the empirical credibility of the primary argument.
@@
As mentioned earlier, there is no definitive ground truth in this case. Accordingly, the evaluation metrics in this case follow the principles outlined in Section 5 and rely on predictability-based measures such as consistency and consensus, rather than on absolute correctness. These metrics operationalise the “no-ground-truth” condition by quantifying relative coherence between human and AI reviewers.
To support the comparative and threshold-based arguments, evidence from Comparative benchmarking was collected through controlled evaluation trials involving 200 synthetic tender cases.
The following statement clearly explains the assurance activities performed, summarises the tangible artefacts produced or collected, and explicitly links those artefacts to the argument they support.
@@
    The differences (marginal risk: e.g., -0.2% gap of inconsistency rate) clearly show that the new system (AI–human) is no worse (slightly better) than the existing system (human-human).
    The differences in decision inconsistency between the AI–human and human–human configurations ($\Delta R=-0.2\%$) show that the new system performs no worse and slightly better than the existing process.
@@
The findings indicate that the AI-based review process can be safely adopted alongside or as a partial replacement for the existing human-human review workflow. Adoption is expected to reduce reviewer workload and resource demand while maintaining consistency, fairness, and transparency in decision-making. Continuous monitoring of agreement and fairness indicators is recommended to ensure that performance remains within the validated acceptance threshold.
@@ 7 Limitations and Future Work
This study aims to establish reusable, composable templates for AI safety cases that reflect the empirical and evolving nature of AI systems. Several limitations remain.
7.1 Scalability of dynamic safety cases
Dynamic safety cases require continuous ingestion, verification, and version tracking of evidence. For continuously evolving systems, achieving through-life safety assurance is challenging due to the fundamental mismatch between rapid, non-deterministic AI evolution and the need for rigorous assurance documentation [Erfan2020dynamic, Shreyas2022dynamic, Laure2025dynamic]. Key limitations include continuous evidence overload from real-time operational data, the computational complexity of automatically updating argument structures when AI behaviour changes, and the lack of unified standards and robust automation infrastructure for managing dynamic evidence and argument revision. Implementing them at the scale of frontier AI systems will require automation infrastructure, tooling standards, and agreement on trusted data pipelines for live safety indicators.
7.2 Interoperability across regulatory regimes
Different jurisdictions emphasise distinct aspects of AI safety, such as capability thresholds, model access controls, or human oversight. Ensuring that reusable templates can be parameterised for diverse regulatory contexts without fragmentation is challenging due to divergent regulatory (e.g., EU AI Act [act2024eu]) philosophies, varied evidential standards and levels of assurance across sectors  [Leveson2011regulation], and the difficulty of translating high-level legal requirements into structured, auditable technical claims within CAE templates [Natsuki2025regulation]. Achieving effective interoperability requires developing modular template components that can be selectively activated and customized for targeted jurisdictions and required assurance levels.
7.3 Evaluation reproducibility
@@
7.4 Relationship to verification processes
This paper treats verification outcomes (e.g., test results, certificates, proofs, or monitoring summaries) as evidence inputs to an AI safety case, but does not model the verification processes that produce them. In particular, we do not characterise how behavioural properties are generated, checked, or enforced at design time or runtime. These concerns are increasingly addressed by verifiability-first AI engineering approaches [zhuverifiability], which focus on structuring systems to enable the verification of behavioural claims at scale. The contribution of this paper is complementary. It focuses on how verified artefacts, once available, are assembled, contextualised, and justified within a structured safety argument.
7.5 Future directions
Further research should operationalise the templates into a tool-supported Safety Case Pattern Library, enabling instantiation and continuous management across AI lifecycles. Key directions include formalising uncertainty propagation across reasoning modes, developing composite safety metrics for dynamic systems, and standardising comparator-based marginal-risk evaluations, all of which represent promising steps for both scientific and regulatory progress. In addition, three priorities emerge for our next step work: First, using LLMs to automatically generate AI safety cases or CAE templates could significantly support dynamic safety case construction. Recent studies (e.g., S14,S51,S53,S74) show that LLMs can generate the structural components of safety case elements (e.g., CAEs) and accelerate early safety documentation, though they struggle to achieve high accuracy and provide traceable evidence without human review. This points to a hybrid human-AI assurance paradigm in which LLMs serve as drafting accelerators, while humans validate the results. Second, integrating AI risk management and assessment frameworks [NOPSEMA2025safetycase] and operational data, (e.g., guardrail logs [shamsujjoha_ICSA-2025], evaluation results [balesni2024evaluations], and AgentOps observability data [dong2024agentops2]) across AI safety case pipeline, could provide a more comprehensive evidence repository to support the full lifecycle of AI safety case construction, verification, and governance (as illustrated in Figure 3 ecosystem pipeline). Third, developing systematic methods for mapping broad legal requirements into specific, measurable technical claims using CAE templates (e.g.,  [Natsuki2025regulation]) would bridge the gap between regulatory mandates and technical assurance. These directions represent critical steps toward making rigorous, evidence-based safety assurance practical and scalable for frontier AI systems.
@@ 8 Conclusion
In this study, we presented a reusable safety-case template structure for AI systems. It uses a clear Claims-Arguments-Evidence (CAE) approach. It also provides AI-specific taxonomies for claim types, argument forms, and evidence families. These elements help practitioners build safety arguments in a consistent and reviewable way. Our proposed templates address recurring assurance issues in many AI deployments. We include end-to-end patterns for evaluation without ground truth, managing dynamic model updates, and making threshold-based risk decisions. This extends safety-case practice to areas where traditional safety cases often do not fit well. It also supports safety cases that can be checked and updated as the system changes.
In addition, we embedded the templates into a continuous assurance pipeline. This pipeline links safety claims to live metrics and governance artifacts for ongoing monitoring. We applied the approach in a real-world case study on a government AI-based tender evaluation system. The case study shows how the templates can be used in practice and what evidence can be collected to support key claims. Overall, this template-based approach provides a reusable foundation for AI safety cases that must evolve over time. However, it is not a complete solution. Key remaining challenges include scaling dynamic safety cases, supporting different regulatory settings, and improving the reproducibility of evaluations.